OpenStudy (anonymous):

The formula for degrees Celsius is: C = 9 over 5 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 2 degrees Celsius is.

OpenStudy (anonymous):

ok C=9/5(F-32) goes to 5/9C=F-32 then 5/9C+32=F

OpenStudy (anonymous):

its just reverse pemdas

OpenStudy (hba):

There is a point where Fahrenheit and Celsius become equal. Do you know what the point is :P